# Uncertainty Makes It Stable: Curiosity-Driven Quantized Mixture-of-Experts

### v1

@@ 6 Discussion
Our findings challenge conventional wisdom in quantized MoE: while heterogeneity boosts adaptability (e.g., 1-2% F1 gains on Quinn), the primary value lies in stability, not raw accuracy-curiosity routing’s 82% variance reduction enables deployment in timing-sensitive applications like autonomous vehicles or hearing aids. Hardware limitations (e.g., Apple M3’s INT8 overhead) explain modest speedups, suggesting future work on ASIC-optimized quantizers. Cross-dataset robustness highlights ESC-50’s sensitivity to diverse sounds, while UrbanSound8K’s tolerance supports aggressive quantization for urban monitoring.
Limitations include single-device (CPU) evaluation; multi-platform testing (e.g., GPU/edge TPUs) could reveal broader gains. We focus on classification - extending to regression or generation tasks remains open. Ethically, our low-carbon focus promotes sustainable AI, but deployment in surveillance raises privacy concerns, mitigated by on-device processing.
@@ 7 Conclusion
Our curiosity-driven quantized Mixture-of-Experts advances efficient edge AI by fusing epistemic uncertainty-based routing with heterogeneous quantization, delivering stable, low-carbon models without sacrificing accuracy. By prioritizing epistemic exploration, we pave the way for predictable quantized systems in resource-constrained environments, with practical guidelines for deployment. Future extensions to multimodal tasks could further enhance real-world applicability.
@@ Appendix A Carbon Emissions Analysis
Single models (1.2M params) emit 22.8 to 55.0 g CO2 during training, while MoE architectures (2.6 to 5.0M params) produce 77.6 to 230.6 g. However, deployment emissions dominate at scale. For 1M daily inferences over 1 year (grid: 367 g CO2/kWh), FP32-Base training emits 22.8 g but deployment produces 2.4 g/year, a 105,000 increase. Q4-Base shows similar patterns: 52.3 g training versus 4.6 g/year deployment. This demonstrates that deployment emissions dominate by approximately 10,000 for models serving more than 1,000 inferences.
@@ B.1 Optimal Configuration
Q4-Base represents the optimal deployment point with 7.03 compression (687 KB) and 0.858 F1 score, within 0.1% of 16-bit performance while consuming 41% less energy than 8-bit. This balance makes it suitable for most edge deployment scenarios.

### v3

@@ 6 Discussion
Our findings demonstrate that the primary value of curiosity-driven MoE lies in stability and adaptive precision allocation, not raw accuracy. Simple Q4-Base achieves 0.858 F1, outperforming all MoE configurations on aggregate accuracy. However, curiosity routing provides capabilities unavailable with fixed quantization: 50–94% F1 variance reduction on ESC-50/Quinn () ensures predictable performance, while the Q8 expert automatically receives uncertain samples (20% lower confidence, ), providing per-sample precision adaptation. UrbanSound8K shows no variance benefit (baseline std=0.007), confirming the mechanism targets genuine epistemic uncertainty. For safety-critical applications such as hearing aids or autonomous vehicle perception, this interpretable, uncertainty-aware routing is essential—knowing which expert processed each input and why matters more than marginal accuracy gains.
The counter-intuitive energy pattern (8-bit consuming 2.7 more energy than 16-bit) reflects ARM’s NEON SIMD instruction set, which natively optimizes 16/32-bit operations. 8-bit inference requires additional pack/unpack overhead absent from native precision paths, a platform-specific effect that would differ on hardware with dedicated INT8 accelerators (e.g., edge TPUs, NVIDIA Tensor Cores). Cross-dataset robustness highlights ESC-50’s sensitivity to diverse sounds, while UrbanSound8K’s tolerance supports aggressive quantization for urban monitoring.
Limitations include single-device (Apple M3 CPU) evaluation; multi-platform testing on GPUs, edge TPUs, and other ARM variants could reveal broader gains and different energy profiles. Our evaluation focuses on audio classification using ESC-50 and UrbanSound8K—standard benchmarks appropriate for edge deployment validation. The uncertainty-based routing mechanism (Eq. 8) is domain-agnostic, operating on routing probabilities rather than audio-specific features. Since spectrograms are processed as images, the framework naturally extends to vision tasks; we commit to discussing extensions to other modalities and larger-scale models. Ethically, our low-carbon focus promotes sustainable AI, but deployment in surveillance raises privacy concerns, mitigated by on-device processing.
@@ 7 Conclusion
We presented a curiosity-driven quantized MoE framework achieving 50–94% F1 variance reduction on challenging datasets and adaptive precision allocation (). While 4-bit quantization remains optimal for raw accuracy (0.858 F1), curiosity routing provides interpretable, uncertainty-aware expert selection critical for safety-sensitive deployments. Variance reduction scales with baseline uncertainty, confirming the mechanism targets genuine epistemic complexity. Future work includes multi-platform evaluation and extension to vision and multimodal tasks.
@@ Appendix A Carbon Emissions Analysis
Single models (1.2M params) emit 22.8 to 55.0 g CO2 during training, while MoE architectures (2.6 to 5.0M params) produce 77.6 to 230.6 g. However, deployment emissions dominate at scale. For 1M daily inferences over 1 year (grid: 367 g CO2/kWh), FP32-Base training emits 22.8 g but deployment produces 2.4 g/year, a 105,000 increase. Q4-Base shows similar patterns: 52.3 g training versus 4.6 g/year deployment. This demonstrates that deployment emissions dominate by approximately 10,000 for models serving more than 1M inferences.
@@ B.1 Optimal Configuration
Q4-Base represents the optimal deployment point with 7.03 compression (687 KB) and 0.858 F1 score, within 0.1% of 16-bit performance while consuming 31% less energy than 8-bit. This balance makes it suitable for most edge deployment scenarios.